OpenStudy (anonymous):

Use linear approximation, i.e. the tangent line, to approximate 15.3^2 as follows: Let f(x)=x^2 and find the equation of the tangent line to f(x) at x=15. Using this, find your approximation for 15.3^2

OpenStudy (anonymous):

The linear approximation of \(f(x)\) at \(x=15\) is given by the line \[g(x)=f(15)+f'(15)(x-15),\] where \(f(x)=x^2\). So, the approximate value of \(15.3^2\) is given by \(g(15.3)\approx f(15.3)\).

OpenStudy (anonymous):

start with f(x)

OpenStudy (anonymous):

a linear approximation means, for a small increment in x, estimate the change in y to a degree of "1" i.e., tangent line at that point so, our point of function is at x="15"

OpenStudy (anonymous):

\[f(x)=x^2\implies f(15)=225\] now, for a small change in 'x', (dx), find change in f (f') \[dy=2xdx\implies \Delta y=2x\Delta x\] and near x=15, \[\Delta y=2\times15\times\Delta x=13\Delta x\] now, our number is "15.3", which differs by 0.3 (=dx) from 15(=x) \[\Delta y=30(0.3)=9\\ f(x+\Delta x)=y+\Delta y\\ f(15.3)=f(15)+\Delta y=225+9=234 \]

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

welcome